# Continuous Domain Generalization

### v1

@@ 6 Conclusion
This paper presents a unified framework for continuous domain generalization, which aligns structured model parameterization with domain-wise continuous variation. We identify the geometric and algebraic foundations for continuous model evolution and instantiate them through a neural Lie transport operator, which enforces structured parameter transitions over the variation space. Beyond idealized settings, we further address realistic imperfections in the descriptor space by introducing a gating mechanism to suppress noise, and a local chart construction to handle partial information. We provide both theoretical analysis and extensive empirical validation on synthetic and real-world datasets, demonstrating strong generalization performance, robustness, and scalability. Future directions include automatic descriptor construction, integration with generative models for latent descriptor inference, and extension to higher-order model structures beyond parameter fields.
@@ A.1 Dataset Details
Figure 5: Train and test set distributions of the 2-Moons dataset.
@@ A.2 Comparison Methods
  • DNA (Nearest Domain Adaptation): This method first trains a global model on all training domains, then fine-tunes it individually for each domain. For test-time prediction, it selects the closest training domain in descriptor space and directly uses its fine-tuned model.
@@ A.3 Model Configuration
Our model comprises four jointly trained components: a per-domain predictive model, a parameter encoder-decoder pair, a neural transport operator, and a gating mechanism. We adopt a task-specific predictive architecture (detailed below) shared across all domains. Each training domain is associated with its own parameter vector, while maintaining structural consistency. The parameters encoder and decoder are four-layer MLPs with ReLU activations and a 32-dimensional latent space. The gating mechanism consists of a linear transformation and a learnable mask. The neural transport operator is implemented as an exponentiated vector field over the descriptor space. It includes two sub-networks: a field network that maps the source descriptor $z_{i}$ to a set of $n$ basis matrices, and a coefficient network that maps the descriptor difference $\Delta z=z_{j}-z_{i}$ to $n$ scalar weights. The resulting transformation is computed by sequentially applying matrix exponentials of the scaled basis matrices as described in Eq. 6. For models with large parameter sizes, we introduce a shared feature extractor and infer only the remaining domain-specific parameters. All experiments are conducted on a 64-bit machine with two 20-core Intel Xeon Silver 4210R CPUs @ 2.40GHz, 378GB memory, and four NVIDIA GeForce RTX 3090 GPUs. Results are averaged over five runs with different random seeds. We use the Adam optimizer and train for 500 epochs, and we specify the architecture as well as other details for each dataset as follows:
@@
    2-Moons The predictive model is a three-layer MLP with 50 hidden units per layer and ReLU activations. The encoder and decoder are both four-layer MLPs with layer dimensions $[1024,512,128,32]$. The transport operator consists of a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$.
@@
    MNIST The shared feature extractor is a convolutional backbone composed of three convolutional layers with channels $[32,32,64]$, each followed by a ReLU activation and a max pooling layer with kernel size 2. The resulting features are flattened and passed through a dropout layer. The per-domain predictive model is a two-layer MLP with a hidden dimension of 128 and output dimension of 10, with dropout rate 0.7. The encoder and decoder are four-layer MLPs with layer dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@
    fMoW The shared feature extractor is a convolutional backbone composed of three convolutional layers with channels $[32,32,64]$, each followed by a ReLU activation and a max pooling layer. Strided convolutions are used for downsampling. The resulting features are flattened and passed through a dropout layer. The per-domain predictive model is a three-layer MLP with hidden dimensions $[128,32]$ and output dimension 10, with dropout rate 0.5. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@
    YearBook The shared feature extractor is a convolutional backbone composed of three convolutional layers with channels $[32,32,64]$, each followed by a ReLU activation and a max pooling layer. The output is flattened and passed through a dropout layer. The per-domain predictive model is a three-layer MLP with hidden dimensions $[128,32]$ and output dimension 2, with dropout rate 0.7. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@
    Traffic The predictive model is a three-layer MLP that takes as input a flattened 24-dimensional vector representing 12 historical inflow and outflow pairs, and outputs a 12-dimensional vector corresponding to 6-step future predictions. The hidden dimension is set to 64. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@ A.4 Complexity Analysis
In our framework, each domain is associated with a parameter vector $\theta\in\mathbb{R}^{D}$, which is transformed via a neural autoencoder composed of a MLP. The encoder maps $\theta$ into a low-dimensional latent embedding $e$, which is then updated via a neural transport operator defined over the descriptor space. The updated embedding is decoded back into $\theta$ to generate a domain-specific model.
The overall complexity of this transformation is $\mathcal{O}(2(Dn+E)+F)$, where $n$ is the width of the first encoder layer, $E$ is the total number of parameters in the remaining layers of the encoder and decoder, and $F$ denotes the parameters in the transport and gating modules. Specifically, $F$ includes the field network, which maps the source descriptor to a small set of basis matrices in the latent space, and the coefficient network, which maps descriptor differences to scalar weights over those bases. Since both networks operate in a low-dimensional space (i.e., same dimensional as $e$), their parameter sizes remain small. The gating module is implemented as a lightweight mask vector over descriptor dimensions and adds negligible overhead. In practice, $D$ is also small because most model parameters, such as those in convolutional or transformer-based feature extractors, are shared across domains. Domain specific parameters is usually applied only to a few final layers. Consequently, the overhead introduced by our framework remains controlled.
@@
Ablation 2-Moons fMoW Traffic
$\mathcal{T}-Plain$ 13.3 $\pm$ 0.4 42.1 $\pm$ 1.1 16.5 $\pm$ 1.0
$\mathcal{T}-noLie$ 34.8 $\pm$ 0.5 37.8 $\pm$ 0.5 16.4 $\pm$ 0.3
w/o Gating - 42.8 $\pm$ 1.2 16.3 $\pm$ 0.4
w/o Chart - 36.8 $\pm$ 0.5 16.1 $\pm$ 0.3
NeuralLTO 3.2 $\pm$ 1.2 35.9 $\pm$ 0.5 15.1 $\pm$ 0.1
A.5 Ablation Study
To understand the contribution of each core component in our framework, we conduct a series of ablation studies on two classification datasets: 2-Moons, fMoW, and a regression dataset Traffic. We focus on the importance of structure-aware transport design, gating mechanisms, and local chart-based inference. All variants are trained and evaluated under the same protocol as the main experiments.
We achieve the following variants, results can be find in Table 2. (1) Plain: replaces the structured transport operator with a plain MLP that directly maps $(z_{i},z_{j},e_{i})$ to $e_{j}$, termed as $\mathcal{T}-Plain$. The performance of the $\mathcal{T}-Plain$ variant demonstrates the limitations of removing geometric and algebraic constraints, as it consistently underperforms our full model across all datasets. The lack of structural bias causes the transport function to overfit to training tuples, failing to extrapolate to unseen domains. (2) noLie: bypass the exponential mapping that projects Lie algebra elements to the Lie group. Instead, the descriptor difference $\Delta z$ is first applied to the field network to produce a raw linear operator, which is then applied directly to the source embedding $e_{i}$ to obtain $e_{j}$. $\mathcal{T}-noLie$ variant reduces the transformation to a descriptor-conditioned linear mapping without enforcing any group structure. It leads to severe performance degradation, especially on the 2-Moons dataset, and remains consistently suboptimal across real-world datasets. This highlights the importance of algebraic structure, without which the learned transport fails to align reliably with the underlying domain variation. We further ablate two structural modules designed to handle imperfect descriptors. Disabling the gating mechanism leads to performance drops under noisy or redundant descriptors, indicating its role in suppressing irrelevant dimensions. Removing charting similarly impairs generalization under incomplete descriptors, as transport lacks global consistency. Both components are critical for robust modeling under imperfect descriptor conditions. More detailed evaluations of these modules are presented in Section 5.3.
@@ A.5 Ablation Study
A.6 Scalability Analysis
We evaluate the scalability of our framework with respect to two key factors: the number of predictive model parameters and the number of domains. Computation time is measured under increasing model and domain complexity, normalized by the shortest time to provide a consistent basis.
Fig. 7 (a) reports computation time on the fMoW dataset as a function of predictive model size, varied from 80K to 9M parameters by adjusting network depth and width. The growth in runtime is smooth and nearly linear, suggesting that the encoder-decoder pipeline and transport operator introduce only moderate overhead even as parameter size increases. This aligns with the theoretical complexity described in Section A.4. Fig. 7 (b) presents the runtime as a function of the number of domains, evaluated on the 2-Moons dataset with domain counts ranging from 10 to 1000. Domains are sampled from the descriptor space with a fixed number of data points per domain to ensure comparability. The runtime increases linearly with domain count, reflecting the scalability of our per-domain inference pipeline and supporting its applicability to large-scale domain generalization tasks.
@@ A.8 Convergence Analysis
We evaluate convergence under increasing numbers of training domains on the 2-Moons dataset. Specifically, we vary the number of training domains from 5 to 100, while fixing the test domains on a predefined mesh grid to ensure consistent evaluation. As shown in the left panel of Fig. 9, the test error drops rapidly as the number of training domains increases. With as few as 20 training domains, the model already achieves a test error below 10%, and beyond 50 domains, the error consistently falls below 5%. At 80 domains, the model shows near-perfect generalization. These results demonstrate that our model can recover the underlying parameter field even under very sparse supervision, and increasing the number of training domains is crucial for capturing the structured variation.
@@
The center panel of Fig. 9 illustrates the training loss trajectories over epochs under different numbers of training domains (10,30,50,70). Across all settings, the loss consistently decreases and stabilizes after around 120 epochs, demonstrating the convergence stability of our framework. Moreover, as the number of domains increases, the model not only converges to a lower final loss (insight from the previous paragraph) but also reaches convergence more rapidly. This confirms that greater domain coverage not only improves model accuracy but also accelerates training.
@@
Fig. 9 (right) plots both the training loss and test error under our main experimental setting, with over 600 epochs. The model quickly converges and maintains consistent test performance throughout the entire training period, with no signs of overfitting. This stability arises from the structural constraints imposed by the architecture, which regularize domain-wise variation and prevent overfitting to individual domains. This observation aligns with our theoretical analysis in Property 2, suggesting that enforcing structure constraints over the descriptor space naturally limits over-parameterization and supports robust generalization.
@@ Appendix B Optimization Details
The joint training strategy follows the general optimization paradigm established in work [6] on temporal domain generalization, which has been shown to be principled and empirically validated for coordinating model parameter evolution over time. Our method generalizes this idea to structured transport operator across descriptor space, extending its applicability beyond purely temporal settings. Specifically, we jointly optimize the per-domain parameters ${\theta(z_{i})}$, encoder $f_{e}$, decoder $f_{d}$, neural transport operator $\mathcal{T}$, and gating mechanism. For each training descriptor $z_{i}$, we supervise its neighborhood $\mathcal{N}(z_{i})$ by aligning predictions, latent representations, and transported parameters between $(z_{i},z_{j})$ pairs. This formulation ensures local consistency and smooth transitions across the parameter manifold. During inference, we identify the nearest training descriptor for a test domain and apply the transport operator to infer its model parameters.
@@
Algorithm 1 Training and Inference Procedure of NeuralLTO

### v2

@@ 6 Conclusion
This paper presents a unified framework for continuous domain generalization, which aligns structural model parameterization with domain-wise continuous variation. We identify the geometric and algebraic foundations for continuous model evolution and instantiate them through a neural Lie transport operator, which enforces structure-preserving parameter transitions over the variation space. Beyond idealized settings, we further address realistic imperfections in the descriptor space by introducing a gating mechanism to suppress noise, and a local chart construction to handle partial information. We provide both theoretical analysis and extensive empirical validation on synthetic and real-world datasets, demonstrating strong generalization performance, robustness, and scalability.
Building upon these insights, several directions emerge beyond the current setting. First, many seemingly discrete domain generalization tasks may in fact represent sparse samples of an underlying continuous process governed by latent semantics. Discovering these latent descriptors through unsupervised or generative modeling is appealing. Second, scaling the framework toward more expressive model structures offers a promising avenue for extending its applicability to broader domains. Third, exploring arithmetic operations between models in the functional space may further deepen the understanding of continuous generalization.
@@ A.1 Dataset Details
Figure 5: Train and test domain descriptors for the 2-Moons dataset. The left plot shows the 50 training domains, while the right plot shows the 150 randomly sampled test domains and additional test domains uniformly distributed over a fixed mesh grid in the descriptor space.
@@ A.2 Comparison Methods
  • NDA (Nearest Domain Adaptation): This method first trains a global model on all training domains, then fine-tunes it individually for each domain. For test-time prediction, it selects the closest training domain in descriptor space and directly uses its fine-tuned model.
@@ A.3 Model Configuration
We provide full implementation details and code in our repository [^4] https://github.com/Zekun-Cai/NeuralLio. Our model comprises four jointly trained components: a per-domain predictive model, a parameter encoder-decoder pair, a neural transport operator, and a gating mechanism. We adopt a task-specific predictive architecture (detailed below) shared across all domains. Each training domain is associated with its own parameter vector, while maintaining structural consistency. The parameters encoder and decoder are four-layer MLPs with ReLU activations. The gating mechanism consists of a linear transformation and a learnable mask. The neural transport operator is implemented as an exponentiated vector field over the descriptor space. It includes two sub-networks: a field network that maps the source descriptor $z_{i}$ to a set of basis matrices, and a coefficient network that maps the descriptor difference $\Delta z=z_{j}-z_{i}$ to a set of scalar weights. The resulting transformation is computed by sequentially applying matrix exponentials of the scaled basis matrices as described in Eq. 6. For models with large parameter sizes, we introduce a shared feature extractor and infer only the remaining domain-specific parameters. All experiments are conducted on a 64-bit machine with two 20-core Intel Xeon Silver 4210R CPUs @ 2.40GHz, 378GB memory, and four NVIDIA GeForce RTX 3090 GPUs. Results are averaged over five runs with different random seeds using the Adam optimizer. We specify the task-specific predictive architecture for each dataset as follows:
@@
    2-Moons The predictive model is a three-layer MLP with 50 hidden units per layer and ReLU activations. The encoder and decoder are both four-layer MLPs with layer dimensions $[1024,512,128,32]$. The transport operator consists of a 32-dimensional linear field network with 2 basis matrices. The learning rate is set to $1\times 10^{-3}$.
@@
    MNIST The shared feature extractor is a convolutional backbone composed of three convolutional layers with channels $[32,32,64]$, each followed by a ReLU activation and a max pooling layer with kernel size 2. The resulting features are flattened and passed through a dropout layer. The per-domain predictive model is a two-layer MLP with a hidden dimension of 128 and an output dimension of 10. The encoder and decoder are four-layer MLPs with layer dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 3 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@
    fMoW ResNet-50 backbone pretrained on ImageNet as the shared feature extractor to capture high-level semantics. The extracted features are fed into a per-domain predictive model implemented as a three-layer MLP with hidden dimensions $[128,64]$ and an output dimension of 10. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 128-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$.
@@
    YearBook The shared feature extractor is a convolutional backbone composed of three convolutional layers with channels $[32,32,64]$, each followed by a ReLU activation and a max pooling layer. The output is flattened and passed through a dropout layer. The per-domain predictive model is a three-layer MLP with hidden dimensions $[128,32]$ and output dimension 2. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$.
@@
    Traffic The predictive model is a three-layer MLP that takes as input a flattened 96-dimensional vector representing 48 historical inflow and outflow pairs, and outputs a 6-dimensional vector corresponding to 3-step future predictions. The hidden dimension is set to 64. The encoder and decoder are four-layer MLPs with dimensions $[1024,512,128,32]$. The neural transport operator is implemented as a 32-dimensional linear field network with 5 basis matrices. The learning rate is set to $1\times 10^{-3}$ for all components.
@@ A.4 Complexity Analysis
In our framework, all domains share a common feature extractor, while each domain is associated with a domain-specific parameter vector $\theta\in\mathbb{R}^{D}$. This vector is first encoded into a low-dimensional latent embedding $e$ through a neural autoencoder implemented as an MLP. The embedding $e$ is then updated by a neural transport operator defined over the descriptor space and subsequently decoded back into $\theta$ to produce the generalized model.
The overall complexity of this transformation is $\mathcal{O}(2(Dn+E)+F)$, where $n$ is the width of the first encoder layer, $E$ is the total number of parameters in the remaining layers of the encoder, and $F$ denotes the parameters in the transport and gating modules. Specifically, $F$ includes the field network, which maps the source descriptor to a small set of basis matrices in the latent space, and the coefficient network, which maps descriptor differences to scalar weights over those basis matrices. Since both networks operate in a low-dimensional space (i.e., the same dimension as $e$), their parameter sizes remain small. The gating module is implemented as a lightweight mask vector over descriptor dimensions and adds negligible overhead. In practice, $D$ is also small because most model parameters, such as those in convolutional or transformer-based feature extractors, are shared across domains. Domain-specific parameters are usually applied only to a few final layers. Consequently, the overhead introduced by our framework remains controlled.
A.5 Ablation Study
To understand the contribution of each core component in our framework, we conduct a series of ablation studies on the 2-Moons and Traffic datasets. We focus on the importance of structure-aware transport design, gating mechanisms, and local chart-based inference. All variants are trained and evaluated under the same protocol as the main experiments.
@@ A.5 Ablation Study
Ablation 2-Moons Traffic
$\mathcal{T}-Plain$ 13.3 $\pm$ 0.4 16.5 $\pm$ 1.0
$\mathcal{T}-noLie$ 34.8 $\pm$ 0.5 16.4 $\pm$ 0.3
w/o Gating - 16.3 $\pm$ 0.4
w/o Chart - 16.1 $\pm$ 0.3
NeuralLio 3.2 $\pm$ 1.2 15.1 $\pm$ 0.1
We achieve the following variants, results can be found in Table 3. (1) Plain: replaces the neural Lie transport operator with a plain MLP that directly maps $(z_{i},z_{j},e_{i})$ to $e_{j}$, termed as $\mathcal{T}-Plain$. The performance of the $\mathcal{T}-Plain$ variant demonstrates the limitations of removing geometric and algebraic constraints, as it consistently underperforms our full model across all datasets. The lack of structural bias causes the transport function to overfit to training tuples, failing to extrapolate to unseen domains. (2) noLie: bypass the exponential mapping that projects Lie algebra elements to the Lie Group. Instead, the descriptor difference $\Delta z$ is first applied to the field network to produce a raw linear operator, which is then applied directly to the source embedding $e_{i}$ to obtain $e_{j}$. $\mathcal{T}-noLie$ variant reduces the transformation to a descriptor-conditioned linear mapping without enforcing any group structure. It leads to severe performance degradation, especially on the 2-Moons dataset, and remains suboptimal on the real-world dataset. This highlights the importance of algebraic structure, without which the learned transport fails to align reliably with the underlying domain variation. We further ablate two structural modules designed to handle imperfect descriptors. Disabling the gating mechanism leads to performance drops under noisy or redundant descriptors, indicating its role in suppressing irrelevant dimensions. Removing charting similarly impairs generalization under incomplete descriptors, as transport lacks global consistency. Both components are critical for robust modeling under imperfect descriptor conditions. More detailed evaluations of these modules are presented in Section 5.3.
@@
A.6 Scalability Analysis
We evaluate the scalability of our framework with respect to two key factors: the number of prediction model parameters and the number of domains. Computation time is measured under increasing model and domain complexity, normalized by the shortest time to provide a consistent basis.
Fig. 7 (a) reports computation time on the fMoW dataset as a function of prediction model size, varied from 80K to 9M parameters by adjusting network depth and width. The growth in runtime is smooth and nearly linear, which aligns with the theoretical complexity described in Section A.4. Fig. 7 (b) presents the runtime as a function of the number of domains, evaluated on the 2-Moons dataset with domain counts ranging from 10 to 1000. Domains are sampled from the descriptor space with a fixed number of data points per domain to ensure comparability. The runtime increases linearly with domain count, reflecting the scalability of our per-domain inference pipeline and supporting its applicability to large-scale domain generalization tasks.
@@ A.8 Convergence Analysis
We evaluate convergence under increasing numbers of training domains on the 2-Moons dataset. Specifically, we vary the number of training domains from 5 to 100, while fixing the test domains on a predefined mesh grid to ensure consistent evaluation. As shown in the left panel of Fig. 9, the test error drops rapidly as the number of training domains increases. With as few as 20 training domains, the model already achieves a test error below 10%, and beyond 50 domains, the error consistently falls below 5%. At 80 domains, the model shows near-perfect generalization. These results demonstrate that our model can recover the underlying parameter field even under very sparse supervision, and increasing the number of training domains is crucial for capturing the variation structure.
@@
The center panel of Fig. 9 illustrates the training loss trajectories over epochs under different numbers of training domains (10,30,50,70). Across all settings, the loss consistently decreases and stabilizes after around 120 epochs, demonstrating the convergence stability of our framework. Moreover, as the number of domains increases, the model not only converges to a lower final loss (insight from the previous paragraph) but also reaches convergence more rapidly. This confirms that greater domain coverage not only improves model accuracy but also accelerates optimization.
@@
The right panel of Fig. 9 plots both the training loss and test error under our main experimental setting, with over 600 epochs. The model converges rapidly and maintains consistent test performance throughout training, showing no signs of performance degradation. This stability arises from the structural constraints imposed by the architecture, which regularize domain-wise variation and prevent the model from memorizing individual domains. This observation aligns with our theoretical analysis in Property 2, suggesting that enforcing structural constraints naturally limits model overfitting and supports robust generalization.
@@ Appendix B Optimization Details
The joint training strategy follows the general optimization paradigm established in [6], which has been shown to be principled and empirically validated for coordinating the temporal evolution of model parameters. Our method generalizes its idea to a structural transport operator across descriptor space, extending its applicability beyond purely temporal settings. Specifically, we jointly optimize the per-domain parameters $\theta(z_{i})$}, encoder $f_{e}$, decoder $f_{d}$, neural transport operator $\mathcal{T}$, and gating mechanism. For each training descriptor $z_{i}$, we supervise its neighborhood $\mathcal{N}(z_{i})$ by aligning predictions, latent representations, and transported parameters between $(z_{i},z_{j})$ pairs. This formulation ensures local consistency and smooth transitions across the parameter manifold. During inference, we identify the nearest training descriptor for a test domain and apply the transport operator to infer its model parameters. The detailed training and inference procedure are summarized in Algorithm 1.
@@
Algorithm 1 Training and Inference Procedure of NeuralLio